OpenStudy (bloomlocke367):

What is the equation of a parabola with the given vertex and focus? vertex: (-2,5) focus: (-2,6)

OpenStudy (bloomlocke367):

@iGreen

OpenStudy (bloomlocke367):

I don't understand the focus part. The only thing I know about the focus is it's equidistant to the vertex as the directrix is to the vertex

OpenStudy (bloomlocke367):

I know that vertex form is y=a(x-h)^2+k but how do I find a?

OpenStudy (igreen):

I'm not sure either..

OpenStudy (bloomlocke367):

ohhhh wait I think I know. would I plug in the x and y from the focus into the vertex form to find a?

OpenStudy (bloomlocke367):

@iGreen

OpenStudy (anonymous):

first off, do you know what this looks like? if so, it is real easy if not, it is extra confusing

OpenStudy (bloomlocke367):

no...

OpenStudy (bloomlocke367):

I just have those two coordinates

OpenStudy (anonymous):

yes i know that the question is, given those two points, the vertex and the focus, do you know what the parabola looks like? i.e. does it open left, right, up, or down? we need that first

OpenStudy (bloomlocke367):

I have no clue. I'm assuming it opens up because the focus is one unit above the vertex and the directrix would have to be one unit below... but my reasoning could be wrong :/

OpenStudy (anonymous):



OpenStudy (anonymous):

because the vertex is below the focus, this means it opens up and yes, the directrix is one unit below the vertex, so it is \(y=4\) but we don't need that for the equation, now we are done in one step

OpenStudy (anonymous):



OpenStudy (anonymous):

general form is \[4p(y-k)=(x-h)^2\] where the vertex is \((h,k)\) here \(p=1\) the distance between the focus and the directrix, so \(4p=4\) and \((h,k)=(-2,5)\)

OpenStudy (anonymous):

you see once you know what it looks like, it is real easy plug them in and get \[4(y-5)=(x+2)^2\]

OpenStudy (bloomlocke367):

okay, so then you'd distribute the 4 then just solve for y?

OpenStudy (bloomlocke367):

Oh and do (x+2)(x+2)

OpenStudy (anonymous):

i would leave it like it is, since that is a general form, but if you need to solve for \(y\) go head i would not square it though

OpenStudy (bloomlocke367):

why not?

OpenStudy (anonymous):

\[4y-20=(x+2)^2\\ 4y=(x+2)^2+20\\ y=\frac{1}{4}(x+2)^2+5\]

OpenStudy (anonymous):

it depends entirely on the form you want to put it in

OpenStudy (bloomlocke367):

it didn't specify...

OpenStudy (anonymous):

if you want standard for \\[y=ax^2+bx+c\] then go ahead if you want "vertex form" then the last line above

OpenStudy (anonymous):

i would leave it as \[4(y-5)=(x+2)^2\]and be done with it

OpenStudy (bloomlocke367):

okay. I just thought I needed to go from vertex form to get to standard form... but okay. thanks for your help :)